# Timescales diagnostics for saving viscous and MHD-driven dusty discs from external photoevaporation

### v1

@@ 3.3 Disc lifetime diagnostics
Figure 5: Disc diagnostics as a function of $\psi_{\mathrm{DW}}=\alpha_{\mathrm{DW}}/\alpha_{\mathrm{SS}}$ for different total $\alpha_{\mathrm{tot}}=\alpha_{\mathrm{DW}}+\alpha_{\mathrm{SS}}$ (different colours and line styles, see legends in the last column), and for different FUV fluxes over different rows (see label above each row). In all cases, the disc’s initial size ${R_{\mathrm{c}}}=100\,\mathrm{AU}$. Upward pointing arrows at $t=31.6\,\mathrm{My}$ show that the data points represent lower bounds. Left column: $t_{\mathrm{dep},\mathrm{dust}}$, the time when the total dust mass drops below 1/1000 of the initial dust mass. Second column: $t_{\mathrm{dep},\mathrm{gas}}$, the time when the gas mass drops below 1/1000 of the initial gas mass. Third column: $t_{\mathrm{acc.}}$, the time at which the accretion rate onto the star drops below $10^{-11}\,M_{\odot}/\mathrm{yr}$. Fourth column: $t_{\mathrm{IR.ex}}$, the time when the disc would be undetectable due to lack of IR excess; we estimate this as the time when the surface density of small dust at $r=1\,\mathrm{AU}$ drops below $1/\kappa_{0}$, with a reference opacity of small dust grains $\kappa_{0}=1000\,\mathrm{cm}^{2}/\mathrm{g}$. Last column: fraction of dust mass lost to wind (note the somewhat inconsistent vertical axis range due to values spanning different orders of magnitude across the different setups; see also Fig. 4). As shown in the legend in the first and fourth columns (which relate to dust timescales), the upward triangles indicate runs where the dust disc radii initially expand due to significant diffusion, while circles indicate those setups which do not (cfr. top row of Fig. 3).
@@
The results for these diagnostics are shown in Figures 5 and 6 for various $R_{\mathrm{c}}$} and FUV fluxes, as a function of $\psi_{\mathrm{DW}}$ and $\alpha_{\mathrm{tot}}$ (see also Figs. 10, 11, and 12). As we would expect, all timescales are shorter with higher external FUV fluxes, with a clear monotonic dependence (except for small $R_{\mathrm{c}}$} and high $\psi_{\mathrm{DW}}$, see below). The short lifetimes of discs under high FUV fluxes coupled with high accretion rates is a known problem (so-called ‘proplyd lifetime problem’ due to the high mass-loss rates observed, e.g. Henney and O’Dell 1999, Störzer and Hollenbach 1999, O’Dell 2001, and more recently e.g. Winter et al. 2019). This raises the question of whether the angular momentum transport mechanism within the disc plays a role, and naively one would expect that DW discs, since they don’t spread out, would be less susceptible to the effects of external FUV irradiation and thus possibly live longer.
In fact, we typically find that the opposite is true. Because the driving factor to dust evolution is radial drift, the main difference does come from the ability of the disc to spread out due to diffusion. Like in Figure 3, we indicate those discs whose dust radii are able to initially expand with upward triangles in the first and fourth columns of Figures 5, 6 while those that do not are labeled with circles. Strikingly, however, this implies that dust in viscous discs actually tends to live longer than in purely DW disc. Indeed, we find that, although viscosity can replenish the outer regions thus making VE discs more prone to losing dust via FUV winds (Fig. 3, middle row), this effect is completely overridden by the overall efficacy of dust radial drift driving accretion onto the central star (via the combined effect of the various processes mentioned at the beginning of Sect. 3): the dust in non radially-expanding discs has less of a radial distance to cover, and is thus removed more quickly. We therefore conclude that MHD winds do not protect the dust component of smooth discs from being eroded by moderate-to-strong external FUV environments, and in fact the opposite is true.
@@
Results for $t_{\mathrm{IR.ex}}$. Concerning $t_{\mathrm{IR.ex}}$, we see that trends with $\psi_{\mathrm{DW}}$ are similar to $t_{\mathrm{dep},\mathrm{dust}}$: larger $\psi_{\mathrm{DW}}$ lead to shorter $t_{\mathrm{IR.ex}}$, with this difference being especially noticeable for those discs that are able to spread outward. Because $t_{\mathrm{IR.ex}}$ tracks small dust, and most of the mass is in the large grains, we typically have that $t_{\mathrm{IR.ex}}<t_{\mathrm{dep},\mathrm{dust}}$. We also note that the dependence on $\alpha_{\mathrm{tot}}$ is less straightforward for $t_{\mathrm{IR.ex}}$ than in $t_{\mathrm{dep},\mathrm{dust}}$. While $t_{\mathrm{dep},\mathrm{dust}}$ shows a rather clear monotonicity with $\alpha_{\mathrm{tot}}$ (for a given fixed $\psi_{\mathrm{DW}}$), with higher $\alpha_{\mathrm{tot}}$ typically leading to a smaller $t_{\mathrm{dep},\mathrm{dust}}$ (except again for vanishing $\psi_{\mathrm{DW}}$ and small $R_{\mathrm{c}}$} as noted above, again due to disc spreading) this is not so for $t_{\mathrm{IR.ex}}$, which is typically lowest for intermediate $\alpha_{\mathrm{tot}}$. The only exception to this is found for initially small discs and for $\psi_{\mathrm{DW}}\gtrsim 10$.
@@
Results for $f_{\mathrm{PE.wind}}$. Finally, looking at the dust mass fraction removed from the external photoevaporative wind, we find as expected (cfr. Fig. 4 and subsect. 3.2) that $f_{\mathrm{PE.wind}}$ is higher for higher $\alpha_{\mathrm{SS}}$ but it is insensitive to $\alpha_{\mathrm{DW}}$ since wind-driven advection does not replenish the outer regions with dust material (as one can most easily see from the limiting cases $\psi_{\mathrm{DW}}\ll 1$ and $\psi_{\mathrm{DW}}\gg 1$). For low total $\alpha_{\mathrm{tot}}$, this dependence with $\psi_{\mathrm{DW}}$ is less severe, as even pure VE discs with low levels of turbulence have a hard time feeding dust into the outer regions. Moreover, at fixed $\psi_{\mathrm{DW}}\lesssim 10$, the increase of $f_{\mathrm{PE.wind}}$ with $\alpha_{\mathrm{SS}}$ is somewhat lessened for initially large discs once $\alpha_{\mathrm{SS}}\lesssim 10^{-3}$, since the disc is already large enough and photoevaporation efficiencies will be rather similar. Indeed we also find that, as expected, higher FUV fluxes will lead to larger $f_{\mathrm{PE.wind}}$ (please note the different vertical scales in the last columns of Figs. 5 and 6 to accommodate for very different $f_{\mathrm{PE.wind}}$ ranges across all setups). For a fixed $\alpha_{\mathrm{tot}}$, $f_{\mathrm{PE.wind}}$ typically decreases with $\psi_{\mathrm{DW}}$, until it becomes almost insensitive to it once $\psi_{\mathrm{DW}}\gtrsim 10$ (except in one case, namely ${R_{\mathrm{c}}}=100\,\mathrm{AU}$, $\mathrm{FUV}=10\,G_{0}$ where this trend is reversed, albeit by an extremely small amount). We stress again that, in this work, $f_{\mathrm{PE.wind}}$ only tracks the dust that is removed by the external FUV wind, and we are not tracking dust that may be removed by MHD-disc winds (cfr. Appendix B).
@@ 4 Discussion
One of the main findings of our work is that, unlike one may naively expect, the angular momentum transport mechanism plays a rather limited role in modulating the lifetimes of dusty discs exposed to higher ($\gtrsim 100\,G_{0}$) FUV fluxes, and, if anything, DW discs tend in fact to live less than VE discs, both in the gas and in the dust. While the lifetime of the gaseous component still depends on the total $\alpha_{\mathrm{tot}}=\alpha_{\mathrm{SS}}+\alpha_{\mathrm{DW}}$, the dust lifetimes are instead less sensitive to it. In any case, as one would expect, lower $\alpha_{\mathrm{tot}}$ lead to a slower evolution overall, and thus somewhat longer disc lifetimes. As we showed in the previous section, this is mostly due to radial drift of dust, which falls onto the star rather quickly. The role of radial drift was already pointed out by Sellek et al. (2020) in the purely viscous case, who showed that it mitigates the external photoevaporation’s role in raising the dust to gas ratio in the disc. Since higher $\psi_{\mathrm{DW}}$ do not help, one may turn to $\alpha_{\mathrm{tot}}$. In principle, small $\alpha_{\mathrm{tot}}$ values do help in keeping dusty discs around for longer, at least for mild FUV fluxes, but as we noted above, the difference is rather small for more highly irradiated discs (FUV flux $\gtrsim 100\,G_{0}$, bottom rows of Figs. 5, 6). This points to the role of processes in the disc that can halt the inward flux of solids, most notable substructures (Whipple 1972, Pinilla et al. 2012). For purely viscous discs, dust traps can be able to extend the dust lifetime for FUV fluxes up to $\lesssim 10^{3}\,G_{0}$ as long as they are located inside the truncation radius (Gárate et al. 2024). Substructures are often observed in nearby discs in low-$G_{0}$ environments (ALMA Partnership et al. 2015, Andrews et al. 2016, 2018, Bae et al. 2023, Vioque et al. 2025). Some discs experiencing moderate-to-high irradiation have also been observed. Huang et al. (2024) found evidence of substructures in 5 discs out of the 8 discs observed in $\sigma$ Orionis (FUV flux $\simeq 10^{2}-10^{3}\,G_{0}$) with the highest ALMA resolution possible. Otter et al. (2021) observed 72 spatially resolved sources in the Orion Nebula, but with lower resolution, with some hints of the presence of substructures, although the bright large-scale emission makes it difficult to characterize disc morphology. Our work suggests that, when demographic surveys are performed for discs in high-$G_{0}$ environments, if we observe old discs with dust, they should also have substructures.
@@ 4.1 Consequences for planet formation
The removal of dust via external winds (large $f_{\mathrm{PE.wind}}$) implies that fewer grains are able to drift inward and replenish the inner regions. This would also have consequences for the chemical abundances (e.g., C/O, overall metallicity, etc.). Although highly irradiated regions are still poorly constrained by observations, recent JWST observations of the high-mass star-forming region NGC 6357 (Ramírez-Tannus et al. 2023, 2025) seem to suggest that discs may be able to retain the ingredients necessary for rocky planet formation even in high FUV environments. Ndugu et al. (2024) underlined the role of efficient radial drift, which may imply that external FUV fluxes even up to $10^{4}\,G_{0}$ do not cause significant differences in chemical abundances of the inner regions compared to non-irradiated discs. However, they did not model dust entrainment in the external photoevaporative wind. More recently, Calahan et al. (2025) used thermochemical modeling (but not including drift) to conclude that inner disc chemistry may be affected, but the impact would be noticeable only for highly irradiated discs ($\gtrsim 10^{6}\,G_{0}$). We will extend the current model to include chemical abundances in gas and dust phases in a future paper.
@@ 5 Conclusions
The fraction $f_{\mathrm{PE.wind}}$ of dust mass removed by the photoevaporative wind depends naturally on the FUV flux and the initial disc radius. It also shows a dependence on $\alpha_{\mathrm{DW}}/\alpha_{\mathrm{SS}}$ across all FUV fluxes considered here, at least for discs that are initially small and/or for higher total $\alpha_{\mathrm{tot}}$. As a general rule, $f_{\mathrm{PE.wind}}$ is higher for viscous discs (low $\alpha_{\mathrm{DW}}/\alpha_{\mathrm{SS}}$), as these spread outwards and can replenish the outer disc regions; however, this effect is somewhat less prominent for discs that are initially already large (as they don’t need to rely on diffusion to bring material outwards, and they are mainly eroded from the outside-in by the external photoevaporation), and with total $\alpha_{\mathrm{tot}}\sim 10^{-4}$. This will have consequences for planet formation as a higher $f_{\mathrm{PE.wind}}$ restricts the availability of solids and the chemical inventory in the inner disc.
Future work will address some limitations of this investigation, including additional effects which can give rise to substructures (e.g. internal photoevaporation or growing planetary cores). We will also add the tracing of chemical species to track the volatile delivery in the inner-most regions of discs, make predictions about important observables such as C/O ratios and metallicity.
@@ Appendix A Comparison of photoevaporative models
For honest comparison with previous works which used the original FRIED grid of Haworth et al. (2018) (FRIEDv1), we show the outcome of a few setups with the original grid and with the updated version from Haworth et al. (2023) (FRIEDv2) we have used throughout the paper. The differences in the mass-loss rates arise from uncertainties in the wind microphysics, specifically the PAH-to-gas ratio $f_{\mathrm{PAH}}$ (more conservative in the original FRIED grid), the UV cross section of the wind (larger by about a factor 2 in the FRIEDv2 grid), and the dust-to-gas mass ratio in the wind (lower by a factor 3 in the updated grid), see Sect. 4 of Haworth et al. (2023). The left panel of Figure 7 shows the surface density evolution for purely viscous discs with $\alpha_{\mathrm{SS}}=10^{-3}$, ${R_{\mathrm{c}}}=100\,\mathrm{AU}$ and an FUV flux of $1000\,G_{0}$, using the FRIEDv1 and FRIEDv2 grids. The right panel shows the dust mass fraction removed by the FUV-driven wind for different values of $\alpha_{\mathrm{SS}}$ and FUV fluxes (cfr. Figure 8 of Sellek et al. 2020) with the two different FRIED grids. There are some minor differences in the surface density evolution, specifically in the disc truncation radii (cfr. Subsect. 2.2 for the definition), but only by a factor of about a few. The main difference is in $f_{\mathrm{PE.wind}}$, which can vary by up to one order of magnitude for the lowest FUV fluxes, with the updated FRIED grid yielding (for the fiducial $f_{\mathrm{PAH}}=1$ value) consistently lower $f_{\mathrm{PE.wind}}$. This shows that the outer disc microphysics may have some impact on the dust reservoirs in the inner discs (see also Coleman et al. 2025). While there is no FRIEDv2 analog of the original FRIED grid tables, we investigate different $f_{\mathrm{PAH}}$ values in a separate paper (Pichierri et al. in prep).

### v2

@@ 3.3 Disc lifetime diagnostics
Figure 5: Disc diagnostics as a function of $\psi_{\mathrm{DW}}=\alpha_{\mathrm{DW}}/\alpha_{\mathrm{SS}}$ for different total $\alpha_{\mathrm{tot}}=\alpha_{\mathrm{DW}}+\alpha_{\mathrm{SS}}$ (different colours and line styles, see legends in the last column), and for different FUV fluxes over different rows (see labels on the left along the arrow). In all cases, the disc’s initial size ${R_{\mathrm{c}}}=100\,\mathrm{AU}$. Upward pointing arrows at $t=31.6\,\mathrm{My}$ show that the data points represent lower bounds. Left column: $t_{\mathrm{dep},\mathrm{dust}}$, the time when the total dust mass drops below 1/1000 of the initial dust mass. Second column: $t_{\mathrm{dep},\mathrm{gas}}$, the time when the gas mass drops below 1/1000 of the initial gas mass. Third column: $t_{\mathrm{acc.}}$, the time at which the accretion rate onto the star drops below $10^{-11}\,M_{\odot}/\mathrm{yr}$. Fourth column: $t_{\mathrm{IR.ex}}$, the time when the disc would be undetectable due to lack of IR excess; we estimate this as the time when the surface density of small dust at $r=1\,\mathrm{AU}$ drops below $1/\kappa_{0}$, with a reference opacity of small dust grains $\kappa_{0}=1000\,\mathrm{cm}^{2}/\mathrm{g}$. Last column: fraction of dust mass lost to wind (note the somewhat inconsistent vertical axis range due to values spanning different orders of magnitude across the different setups; see also Fig. 4). As shown in the legend in the first and fourth columns (which relate to dust timescales), the upward triangles indicate runs where the dust disc radii initially expand due to significant diffusion, while circles indicate those setups which do not (cfr. top row of Fig. 3).
@@
The results for these diagnostics are shown in Figures 5 and 6 for various $R_{\mathrm{c}}$} and FUV fluxes, as a function of $\psi_{\mathrm{DW}}=\alpha_{\mathrm{DW}}/\alpha_{\mathrm{SS}}$ and $\alpha_{\mathrm{tot}}$ (see also Figs. 10, 11, and 12). As we would expect, all timescales are shorter with higher external FUV fluxes, with a clear monotonic dependence (except for small $R_{\mathrm{c}}$} and high $\psi_{\mathrm{DW}}$, see below). The short lifetimes of discs under high FUV fluxes coupled with high accretion rates is a known problem (so-called ‘proplyd lifetime problem’ due to the high mass-loss rates observed, e.g. Henney and O’Dell 1999, Störzer and Hollenbach 1999, O’Dell 2001, and more recently e.g. Winter et al. 2019). This raises the question of whether the angular momentum transport mechanism within the disc plays a role, and naively one would expect that DW discs, since they don’t spread out, would be less susceptible to the effects of external FUV irradiation and thus possibly live longer.
In fact, we typically find that the opposite is true. Because the driving factor to dust evolution is radial drift, the main difference does come from the ability of the disc to spread out due to diffusion. Like in Figure 3, we indicate those discs whose dust radii are able to initially expand with upward triangles in the first and fourth columns of Figures 5, 6 while those that do not are labelled with circles. Strikingly, however, this implies that dust in viscous discs actually tends to live longer than in purely DW disc. Indeed, we find that, although viscosity can replenish the outer regions thus making VE discs more prone to losing dust via FUV winds (Fig. 3, middle row), this effect is completely overridden by the overall efficacy of dust radial drift driving accretion onto the central star (via the combined effect of the various processes mentioned at the beginning of Sect. 3): the dust in non radially-expanding discs has less of a radial distance to cover, and is thus removed more quickly. We therefore conclude that MHD winds do not protect the dust component of smooth discs from being eroded by moderate-to-strong external FUV environments, and in fact the opposite is true.
@@
Results for $t_{\mathrm{IR.ex}}$. Concerning $t_{\mathrm{IR.ex}}$, we see that trends with $\psi_{\mathrm{DW}}$ are similar to $t_{\mathrm{dep},\mathrm{dust}}$: larger $\psi_{\mathrm{DW}}$ lead to shorter $t_{\mathrm{IR.ex}}$, with this difference being especially noticeable for those discs that are able to spread outward. Because $t_{\mathrm{IR.ex}}$ tracks small dust, and most of the mass is in the large grains, we typically have that $t_{\mathrm{IR.ex}}<t_{\mathrm{dep},\mathrm{dust}}$. We also note that the dependence on $\alpha_{\mathrm{tot}}$ is less straightforward for $t_{\mathrm{IR.ex}}$ than in $t_{\mathrm{dep},\mathrm{dust}}$. While $t_{\mathrm{dep},\mathrm{dust}}$ shows a rather clear monotonicity with $\alpha_{\mathrm{tot}}$ (for a given fixed $\psi_{\mathrm{DW}}$), with higher $\alpha_{\mathrm{tot}}$ typically leading to a smaller $t_{\mathrm{dep},\mathrm{dust}}$ (except again for vanishing $\psi_{\mathrm{DW}}$ – i.e. VE discs – and small $R_{\mathrm{c}}$} as noted above, again due to disc spreading) this is not so for $t_{\mathrm{IR.ex}}$, which is typically lowest for intermediate $\alpha_{\mathrm{tot}}$. The only exception to this is found for initially small discs and for $\psi_{\mathrm{DW}}\gtrsim 10$.
@@
Results for $f_{\mathrm{PE.wind}}$. Finally, looking at the dust mass fraction removed from the external photoevaporative wind, we find as expected (cfr. Fig. 4 and subsect. 3.2) that $f_{\mathrm{PE.wind}}$ is higher for higher $\alpha_{\mathrm{SS}}$ but it is insensitive to $\alpha_{\mathrm{DW}}$ since wind-driven advection does not replenish the outer regions with dust material (as one can most easily see from the limiting cases $\psi_{\mathrm{DW}}\ll 1$, i.e. VE discs, and $\psi_{\mathrm{DW}}\gg 1$, DW discs). For low total $\alpha_{\mathrm{tot}}$, this dependence with $\psi_{\mathrm{DW}}$ is less severe, as even pure VE discs with low levels of turbulence have a hard time feeding dust into the outer regions. Moreover, at fixed $\psi_{\mathrm{DW}}\lesssim 10$, the increase of $f_{\mathrm{PE.wind}}$ with $\alpha_{\mathrm{SS}}$ is somewhat lessened for initially large discs once $\alpha_{\mathrm{SS}}\lesssim 10^{-3}$, since the disc is already large enough and photoevaporation efficiencies will be rather similar. Indeed we also find that, as expected, higher FUV fluxes will lead to larger $f_{\mathrm{PE.wind}}$ (please note the different vertical scales in the last columns of Figs. 5 and 6 to accommodate for very different $f_{\mathrm{PE.wind}}$ ranges across all setups). For a fixed $\alpha_{\mathrm{tot}}$, $f_{\mathrm{PE.wind}}$ typically decreases with $\psi_{\mathrm{DW}}$, until it becomes almost insensitive to it once $\psi_{\mathrm{DW}}\gtrsim 10$ (except in one case, namely ${R_{\mathrm{c}}}=100\,\mathrm{AU}$, $\mathrm{FUV}=10\,G_{0}$ where this trend is reversed, albeit by an extremely small amount). We stress again that, in this work, $f_{\mathrm{PE.wind}}$ only tracks the dust that is removed by the external FUV wind, and we are not tracking dust that may be removed by MHD-disc winds (cfr. Appendix B).
@@ 4 Discussion
One of the main findings of our work is that, unlike one may naively expect, the angular momentum transport mechanism plays a rather limited role in modulating the lifetimes of dusty discs exposed to higher ($\gtrsim 100\,G_{0}$) FUV fluxes, and, if anything, DW discs tend in fact to live less than VE discs, both in the gas and in the dust. While the lifetime of the gaseous component still depends on the total $\alpha_{\mathrm{tot}}=\alpha_{\mathrm{SS}}+\alpha_{\mathrm{DW}}$, the dust lifetimes are instead less sensitive to it. In any case, as one would expect, lower $\alpha_{\mathrm{tot}}$ lead to a slower evolution overall, and thus somewhat longer disc lifetimes. As we showed in the previous section, this is mostly due to radial drift of dust, which falls onto the star rather quickly. The role of radial drift was already pointed out by Sellek et al. (2020) in the purely viscous case, who showed that it mitigates the external photoevaporation’s role in raising the dust to gas ratio in the disc. Since higher $\psi_{\mathrm{DW}}$ values do not help, one may turn to $\alpha_{\mathrm{tot}}$. In principle, small $\alpha_{\mathrm{tot}}$ values do help in keeping dusty discs around for longer, at least for mild FUV fluxes, but as we noted above, the difference is rather small for more highly irradiated discs (FUV flux $\gtrsim 100\,G_{0}$, bottom rows of Figs. 5, 6). This points to the role of processes in the disc that can halt the inward flux of solids, most notable substructures (Whipple 1972, Pinilla et al. 2012). For purely viscous discs, dust traps can be able to extend the dust lifetime for FUV fluxes up to $\lesssim 10^{3}\,G_{0}$ as long as they are located inside the truncation radius (Gárate et al. 2024). Substructures are often observed in nearby discs in low-$G_{0}$ environments (ALMA Partnership et al. 2015, Andrews et al. 2016, 2018, Bae et al. 2023, Vioque et al. 2025). Some discs experiencing moderate-to-high irradiation have also been observed. In $\sigma$ Orionis (FUV flux $\simeq 10^{2}-10^{3}\,G_{0}$), Huang et al. (2024) used the highest ALMA resolution possible and recovered substructures in 8 out of 8 discs in visibility modelling, with five of them have substructures unambiguously resolved in the image plane. Otter et al. (2021) observed 72 spatially resolved sources in the Orion Nebula, but with lower resolution, with some hints of the presence of substructures, although the bright large-scale emission makes it difficult to characterise disc morphology. Our work suggests that, when demographic surveys are performed for discs in high-$G_{0}$ environments, if we observe old discs with dust, they should also have substructures.
@@ 4.1 Consequences for planet formation
The removal of dust via external winds (large $f_{\mathrm{PE.wind}}$) implies that fewer grains are able to drift inward and replenish the inner regions. This would also have consequences for the chemical abundances (e.g., C/O, overall metallicity, etc.). Although highly irradiated regions are still poorly constrained by observations, recent JWST observations of the high-mass star-forming region NGC 6357 (Ramírez-Tannus et al. 2023, 2025) seem to suggest that discs may be able to retain the ingredients necessary for rocky planet formation even in high FUV environments. Ndugu et al. (2024) underlined the role of efficient radial drift, which may imply that external FUV fluxes even up to $10^{4}\,G_{0}$ do not cause significant differences in chemical abundances of the inner regions compared to non-irradiated discs. However, they did not model dust entrainment in the external photoevaporative wind. More recently, Calahan et al. (2025) used thermochemical modelling (but not including drift) to conclude that inner disc chemistry may be affected, but the impact would be noticeable only for highly irradiated discs ($\gtrsim 10^{6}\,G_{0}$). We will extend the current model to include chemical abundances in gas and dust phases in a future paper.
@@ 5 Conclusions
The fraction $f_{\mathrm{PE.wind}}$ of dust mass removed by the photoevaporative wind depends naturally on the FUV flux and the initial disc radius. It also shows a dependence on $\alpha_{\mathrm{DW}}/\alpha_{\mathrm{SS}}$ across all FUV fluxes considered here, at least for discs that are initially small and/or for higher total $\alpha_{\mathrm{tot}}$. As a general rule, $f_{\mathrm{PE.wind}}$ is higher for viscous discs (low $\alpha_{\mathrm{DW}}/\alpha_{\mathrm{SS}}$), as these spread outwards and can replenish the outer disc regions. However, this effect is somewhat less prominent for discs that are initially already large, as they don’t need to rely on diffusion to bring material outwards, and with total $\alpha_{\mathrm{tot}}\sim 10^{-4}$. This will have consequences for planet formation as a higher $f_{\mathrm{PE.wind}}$ restricts the availability of solids and the chemical inventory in the inner disc.
Future work will address some limitations of this investigation, including additional effects which can give rise to substructures (e.g. internal photoevaporation or growing planetary cores). We will also add the tracing of chemical species to track the volatile delivery in the inner-most regions of discs and make predictions about important observables such as C/O ratios and metallicity.
@@ Appendix A Comparison of photoevaporative models
For honest comparison with previous works which used the original FRIED grid of Haworth et al. (2018) (FRIEDv1), we show the outcome of a few setups with the original grid and with the updated version from Haworth et al. (2023) (FRIEDv2) we have used throughout the paper. The differences in the mass-loss rates arise from uncertainties in the wind microphysics, specifically the PAH-to-gas ratio $f_{\mathrm{PAH}}$ (more conservative in the original FRIED grid), the UV cross section of the wind (larger by about a factor 2 in the FRIEDv2 grid), and the dust-to-gas mass ratio in the wind (lower by a factor 3 in the updated grid), see Sect. 4 of Haworth et al. (2023). The left panel of Figure 7 shows the surface density evolution for purely viscous discs with $\alpha_{\mathrm{SS}}=10^{-3}$, ${R_{\mathrm{c}}}=100\,\mathrm{AU}$ and an FUV flux of $1000\,G_{0}$, using the FRIEDv1 and FRIEDv2 grids. The right panel shows the dust mass fraction removed by the FUV-driven wind for different values of $\alpha_{\mathrm{SS}}$ and FUV fluxes (cfr. Figure 8 of Sellek et al. 2020) with the two different FRIED grids. There are some minor differences in the surface density evolution, specifically in the disc truncation radii (cfr. Subsect. 2.2 for the definition), but only by a factor of about a few. The main difference is in $f_{\mathrm{PE.wind}}$, which can vary by up to one order of magnitude for the lowest FUV fluxes, with the updated FRIED grid yielding (for the fiducial $f_{\mathrm{PAH}}=1$ value) consistently lower $f_{\mathrm{PE.wind}}$. This shows that the outer disc microphysics may have some impact on the dust reservoirs in the inner discs (see also Coleman et al. 2025). While there is no FRIEDv2 analogue of the original FRIED grid tables, we investigate different $f_{\mathrm{PAH}}$ values in a separate paper (Pichierri et al. in prep).